Cash Discount

Go For Broke Mining was extended credit terms of 3/15 net 30 EOM.

The cost of giving up the cash discount, assuming payment would be made on the last day of the credit period, would be ________. If the firm were able to stretch its accounts payable to 60 days without damaging its credit rating, the cost of giving up the cash discount would only be ________.
Answer #1

Cost of giving up the cash discount = (Disc.%/100 - Disc.%) × [365 / (Days Credit Outstanding – Discount Period)]

Cost of giving up the cash discount = (3/97) × [365 / (30 - 15)] = 75.26%

If the firm were able to stretch its accounts payable, the cost of giving up the cash discount would only be:

Cost of giving up the cash discount = (3/97) × [365 / (60 - 15)] = 25.09%